A metallic sphere of radius R is charged by assembling N electrons (electronic charge e) on it. The total work done is..
 

Grade:12th pass

1 Answers

Susmita
5 years ago
It is only the potential V=(eN)/4\pi \varepsilonR
You can calculate it by integrating -E.dr from infinity to R.
E is electric field.
E=(eN)/4\pi \varepsilonr2